2 cars leave the same corner at the same time, traveling at right angles to each other. One car is traveling 3 km per hour faster than the other car. After 1 hour they are 15 km apart. How fast is each car traveling?
Answer by mananth(16949)

let speed of one car be x kmph
other car = x+3 kmph
...
distance traveled by first car in1 hour = x km
distance traveled by other car = x+3 km
..
they are at distance of 15 km
..
apply pythagoras theorem
..
x^2+(x+3)^=15^2
x^2+x^2+6x+9=225
2x^2+6x-216=0
/2
x^2+3x-108=0
x^2+12x-9x-108=0
x(x+12)-9(x+12)=0
(x+12)(x-9)=0
x=9 km /h speed of one car.
speed of other car = 9+3 = 12 km/h
...
m.ananth@hotmail.ca